Please answer with mathematical formulas, not an excel sheet please 3. John borrows an amount at an annual interest rate of 8%. He repays all interest and principal in a lump sum at the end of
Answer #1

Suppose John borrows an amount X and X is also the price of the bond

The bond pays an amount of 50 every 6 months for 5 years (10 payments) and 1000 at the end of 5 years. The discount rate for a 6 month period is 4.5%.  So

X= 50/1.045 + 50/1.0452+ 50/1.0453+...... 50/1.04510 + 1000/1.04510

=50/0.045 * (1- (1/1.045)10) + 1000/1.04510

= 395.636 + 643.928 = 1039.564

So, the final amount that John has to pay after 10 years is 1039.564 * 1.0810 = 2244.34

John reinvests the bond coupons at 5.8% converted semiannually, so for every 6 month,he gets a rate of 2.9% .  After 5 years, the total amount that John will get is given by

50* 1.0299 + 50* 1.0298 + .... + 50 + 1000 = 50 * (1.02910-1) / 0.029 +1000 = 1570.561

This amount is invested at a rate of 6.5 % annually for the remaining 5 years , so this amount will turn out to become

1570.561 * (1.065)5 = 2151.805 after 10 years

The additional amount of 50 which is deposited at the end of year 6 through 10 will amount to the following  

50* 1.0654+ 50* 1.0653+ 50* 1.0652+ 50* 1.0651+50 = 50/0.065* (1.0655-1) = 284.682

In total, John will have in his account 2151.805+284.682 = 2436.487

Out of which , after paying 2244.34 , the balance of 192.147 at the end of 10 years

So, John will have an amount of 192.147 remaining at the end of 10 years